A wire M of diameter d is joint to another wire N of the same material. The diameter of N is half that of M but the two wires have the same length, If a steady potential difference is applied across the ends of the composite wire, which of the following statements is correct?

A. The current in N is half of the current in M

B. The drift velocity of the electrons in N is four times that of the electrons in M.

C. The number of charge carriers per unit length in M is twice that in N

D. The drift velocity of the electrons in M is twice that of the electrons in N


1
Expert's answer
2021-11-04T16:20:49-0400

drift velocity of the electrons:


"v = \frac{I}{enS}"


"I_N=I_M=I(\text{since the serial connection of the conductors})"


"v_N= \frac{I}{enS_N}"


"v_M= \frac{I}{enS_M}"


"\frac{v_N}{v_M}= \frac{S_M}{S_N};"


"\frac{S_M}{S_N}= \frac{\pi\frac{d_M^2}{4}}{\pi\frac{d_N^2}{4}}=4"


"\frac{v_N}{v_M}=4"


"\text{Answer: }"B. The drift velocity of the electrons in N is four times that of the electrons in M.